# Quantum Wave Function Collapse for Procedural Content Generation

### v1

@@ 2.1 Quantum Wave Function Collapse (QWFC)
where $\mathcal{C}_{\sigma}(k)$ represents the set of possible partial content instances from section 2.1.1,
$U_{k}^{\mathrm{ctrl}}(C^{\prime}):=\bigotimes_{(s,v)\in L_{k}(C^{\prime})}\ket{v-1}\bra{v-1}_{s}\in\bigotimes_{l=1}^{k-1}\mathcal{H}_{\sigma_{l}}$ (43)
denotes the control operator, and
$U_{k}^{\mathrm{init}}(C^{\prime}):=\ket{C^{\prime}}\bra{0}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ (44)
the initialization operator with
|$\ket{C^{\prime}}_{\sigma_{k}}:=\sum_{\nu\in A}\sqrt{p_{k}^{\mathrm{v,P}}(\nu|\sigma_{k},C^{\prime})}\ket{\nu-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}},$ (45)
which is based on the pattern-based probability distribution of values $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ from section 1.2.1. The summation in eq. 43 runs over all adjacent segments
$L_{k}(C^{\prime}):=\{(s,v)\,|\,$ $(s,v)\in C^{\prime}\,\land\,\exists\,P\in R_{P}$
$\exists\,d\in P_{d}:\alpha^{d}_{\sigma_{k}s}=1\}\subseteq C^{\prime},$ (46)
where $R_{P}:=\{P_{1},\dots,P_{m}\}$ stands for the set of patterns in the pattern-based ruleset $R\in\mathcal{R}^{\mathrm{P}}$ with $P_{i}\in\mathcal{P}$ for $i\in[1,m]$ and $P_{d}:=\{d_{1},\dots,d_{n}\}$ stands for the set of directions in the pattern $P\in\mathcal{P}$, eq. 21, with $d_{j}\in[1,D]$ for $j\in[1,n]$. In particular, as discussed above for the classic case, choosing $C^{\prime}$ instead of $L_{k}(C^{\prime})$ in the summation has no effect on $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ and thus |$\ket{C^{\prime}}_{\sigma_{k}}$.
In iteration $k$, only the qubits in $\mathcal{Q}_{k-1}^{+}$ with
$\mathcal{Q}_{k}^{+}:=\{q\,|\,q=q_{k}^{i}\in\mathcal{Q}_{\sigma_{k}}\,\forall\,i\in[1,q]\,\land\,\sigma_{k}\in L_{k}^{+}\}$ (47)
are affected by $U^{\mathrm{ctrl}}(C^{\prime})$ for some $C^{\prime}\in\mathcal{C}_{\sigma}(k)$, where
$L_{k}^{+}:=\{s\,|\,\exists\,C^{\prime}\in\mathcal{C}_{\sigma}(k)\,:\,s\in S^{+}(L_{k}(C^{\prime}))\}$ (48)
denotes the affected segments for $k=[1,N]$. Here, $\mathcal{Q}_{k}^{+}\subseteq\{\mathcal{Q}_{1},\dots,\mathcal{Q}_{k}\}$ and $L_{k}^{+}\subseteq\{\sigma_{1},\dots,\sigma_{k}\}$, respectively.
The second operator in eq. 41 is the neutral operator
@@
Initially, all qubits $\mathcal{Q}_{k}$ are in the ground state. Therefore,
$U_{k}^{\mathrm{init},0}:=\sum_{v\in[2,W]}\ket{v-1}\bra{v-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ (50)
vanishes. Furthermore,
@@
$U_{k}^{\mathrm{ctrl},0}(C^{\prime})\!:=\!\left[\bigotimes_{l=1}^{k-1}\mathbbm{1}_{\sigma_{l}}-U_{k}^{\mathrm{ctrl}}(C^{\prime})\right]\otimes\mathbbm{1}_{\sigma_{k}}\!\in\bigotimes_{l=1}^{k}\mathcal{H}_{\sigma_{l}}$ (51)
does not change the state by construction.
@@
The operator $U_{k}$, eq. 41, can be straightforwardly realized with quantum gates for a given $k\in[1,N]$. To this end, only the conditional initialization operator $U_{k}^{\mathrm{coin}}$, eq. 42, has to be considered because all other parts of $U_{k}$ have no effect on the resulting state. The conditional initialization operator consist of a sequence of joint pairs of control-initialization operators $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, one for each $C^{\prime}\in\mathcal{C}_{\sigma}(k)$. Each of these joint pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{\sigma_{k}}$, which is a well-known task that may, however, require an exponential number of gates [11, 12, 13]. A sketch is shown in fig. 4 (b).
For the definition of QWFC, a noise-free, ideal quantum device is presumed, that is able to run the emerging circuits without any errors. However, the currently available noisy intermediate-scale quantum (NISQ) hardware exhibits significant noise and uncertainty [14]. As a consequence, the content sampled from eq. 40 can also be expected to violate the prescribed patterns. This will be demonstrated further below for several PCG use cases.
@@
The proposed conditions S1, S2, V1, and V2 are in fact not mandatory in the original implementation of CWFC, where a violation leads to a restart of the algorithm. This can also be considered in QWFC by adding an additional “conflict detection qubit” that stores this information, an extension that is not further discussed here.
2.2 QWFC Example: Checkerboard
In the following, the above example for the generation of $3\times 3$ checkerboard images is considered for QWFC, where the predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ is used for eq. 34 as visualized in fig. 5 (a). The resulting CWFC circuit is shown in fig. 5 (b) and the resulting probability distribution of content instances $p(C)$, eq. 40, in fig. 5 (c). Given a noise-free quantum device, a measurement of the circuit corresponds to drawing a sample from this distribution.
Figure 5: QWFC for the generation of $3\times 3$ images with black and white checkerboard patterns in analogy to fig. 3. (a) Predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ for eq. 34. (b) Circuit layout with the same symbols as in fig. 4. (c) Generated content instances $C_{1}$ and $C_{2}$ with $p(C_{1})=p(C_{2})=\frac{1}{2}$, eq. 40.
2.3 Hybrid Quantum-Classical WFC (HWFC)
The required number of qubits for the proposed QWFC method increases linearly with the number of segments and logarithmically with the number of symbols in the alphabet as shown in eqs. 36 and 39. At the same time, the respective circuits may become exponentially deeper because of the costly conditional preparation of probability distributions. The reliable execution of such big circuits on NISQ hardware can be very challenging or even impossible.
[…]
Finally, the distribution of the joint content instances is given by
$p\left(C=\bigcup_{h=1}^{H}C^{h}\right)=\prod_{h=1}^{H}p^{h}(C^{h}|C^{1},\dots,C^{h-1}),$ (52)
which corresponds to performing a QWFC on the complete content, cf. fig. 6. With the proposed HWFC method, the size of the QWFC circuits can be reduced at the price of an additional outer iteration loop and the inclusion of classical correlations in the conditional QWFC circuit generation.
3 DEMONSTRATION
In the present section, five simple content creation use cases are presented to demonstrate the proposed QWFC and HWFC methods. For the realization of these examples, both quantum computing simulators and actual IBM Quantum devices were used, which were accessed via the IBM Quantum cloud services [15] during December 2023.
3.1 Checkerboard Revisited
As a first use case, the previously presented checkerboard example is recalled, cf. figs. 2 and 5. Again, $3\times 3$ instances are considered, for which a comparison between QWFC and HWFC is performed on the IBM Quantum device ibm_kyoto (Eagle r3 processor version 1.2.4, 127 qubits). For QWFC, 10 0001000010\,00010 000 shots are measured, each yielding a content instance. For HWFC, $H=3$ equal-sized partitions are used in analogy to fig. 6 and the method is repeated $131$ times, each time yielding a content instance. The results are shown in fig. 7.
[…]
In total, there are $5122^{9}=$ 512 possible content instances $C\in\mathcal{C}$, eq. 1, but only two of them, $C_{1}$ and $C_{2}$, are valid. The remaining $510$ content instances are invalid because they violate the patterns. In fig. 7 (a), QWFC is able to generate the two valid instances, but at the same time also produces a lot of invalid instances due to hardware imperfections. From fig. 7 (b) it can be seen that the results from HWFC are much closer to the idealized solution that would contain only $C_{1}$ and $C_{2}$ with equal probability. It is a consequence of the reduced circuit size from this method that invalid solutions occur with much less frequency than for QWFC.
3.2 Pipes
The second use case is to create an image from tiles of seven different pipe (or line) sections as well as one blank tile. That is, the alphabet consists of $W=8$ symbols and each tile in the image constitutes a segment. The alphabet is shown in fig. 8 (a). As adjacency configuration, the nearest-neighbor adjacency from fig. 1 is presumed with $D=4$. The pattern-based ruleset is chosen in such a way that neighboring tiles have to form a connecting network of pipes. Three example rules for this purpose are shown in fig. 8 (b). In total, $m=2048$ rules are required to achieve this goal.
Figure 8: Pipe generation. The generated image is required to show a network of connected pipes. (a) Alphabet consisting of $W=8$ symbols, each of which represents a tile with a pipe section and an empty tile. (b) Example of three pattern-based rules $r_{1}^{\mathrm{p}}$, $r_{2}^{\mathrm{p}}$, and $r_{3}^{\mathrm{p}}$, cf. fig. 2. (c) Example $10\times 4$ tile images that were generated using HWFC with $H=10$ on the IBM Quantum device ibm_kolkata and on simulator as well as CWFC.
Exemplarily, images consisting of $10\times 4$ tiles ($N=40$) are considered, for which a comparison between HWFC on the IBM Quantum device ibm_kolkata (Falcon r5.11 processor version 1.14.8, 27 qubits), HWFC on a simulator and CWFC is performed. For HWFC, $H=10$ equal-sized partitions are used, each corresponding to one column of the image. A few resulting instances are shown in fig. 8 (c).
3.3 Hexagon Map
The third use case addresses images based on hexagonal tiles that can be interpreted as a map consisting of different terrains. For this purpose, an alphabet of $W=4$ symbols is chosen, each corresponding to a unicolored tile (blue, yellow, green, and gray) and each tile constitutes a segment. As adjacency configuration, a nearest-neighbor adjacency is presumed ($D=6$). The prescribed pattern-based ruleset only allows connections of the form blue-yellow-green-gray, which can be fulfilled with $m=1586$ rules, see fig. 9 (a). To promote a higher occurrence of blue tiles, the factors in eq. 24 are chosen as $u=5$ for blue tiles and $u=1$ otherwise.
As a first example, images consisting of $N=7$ tiles are considered, for which QWFC on a simulator and CWFC are performed. The resulting distributions of content instances $p(C)$ in fig. 9 (b) is exact for QWFC, eq. 40, and based on 10 0001000010\,00010 000 samples for CWFC, eq. 10. Without surprise, both distributions are significantly different. The second example considers images from $N=127$ tiles, for which HWFC with $H=20$ is performed on a simulator. Example instances are shown in fig. 9 (c).
3.4 Platformer
The fourth use case is motivated by computer game level design. The goal is to create a platformer-type level from a set of eight tiles [16], each corresponding to a different game element (ground, grass, mushroom, block, air, and a tree that is composed of three tiles). Consequently, the alphabet consists of $W=8$ symbols, see fig. 10 (a). As adjacency configuration, the two tiles above and below are taken into account as neighbors ($D=2$). A pattern-based ruleset is prescribed that meets five requirements: (i) The bottom row consists of ground tiles and ground tiles can only be placed on top of each other, (ii) only a grass tile or a mushroom tile can be placed above a ground tile, (iii) the tree tiles must be placed in order with the lowest tile above a grass tile, (iv) air tiles can only be placed above grass, mushroom or other air tiles, (v) block tiles can only be placed between two air tiles. These requirements can be fulfilled with $m=15$ rules, see fig. 10 (b). The first requirement can be resolved with a functional factor, eq. 29, that vanishes for all non-ground tiles in the bottom row. To reduce the occurrence of block tiles, the corresponding factor in eq. 24 is chosen as $u=$0.1$$ for $r_{6}^{\mathrm{p}}$. All other factors are set to $u=1$. The segment order $\sigma$ is chosen such that the levels are built up from bottom to top.
Figure 10: Level generation. (a) Alphabet consisting of eight symbols, each of which represents a tile [16]. (b) Pattern-based ruleset consisting of $m=15$ rules. (c) Example $10\times 10$ tile images that were generated using HWFC with $H=20$ on a simulator.
Exemplarily, levels with $10\times 10$ tiles are considered, for which HWFC with $H=20$ equal-sized partitions (each corresponding to a half row of the level) is performed on a simulator. Four resulting instances are shown in fig. 10 (c).
3.5 Voxel Skyline
The final use case considers the creation of a 3d voxel graphic with a binary alphabet of $W=2$ symbols, which represent the presence or absence of a voxel. Each voxel constitutes a segment, as adjacency configuration the voxel above and below are taken into account ($D=2$). The chosen pattern-based ruleset with $m=4$ ensures that voxels are built from the ground up, see fig. 11 (a). This leads to a skyline-like structure.
Figure 11: Skyline use case. (a) Pattern-based ruleset consisting of $m=4$ rules in an analogous representation as in the previous figures. (b) Example images from a $4\times 4\times 4$ voxel grid that were generated using QWFC on the IBM Quantum device ibm_osaka and a simulator. (c) Example images from a $10\times 10\times 5$ voxel grid that were generated using QWFC on a simulator.
Exemplarily, two cases are considered: First, images from a $4\times 4\times 4$ voxel grid, for which QWFC is performed both on a the IBM Quantum device ibm_osaka (Eagle r3 processor version 1.0.3, 127 qubits) and a simulator, see fig. 11 (b). An error violating $r_{3}^{\mathrm{p}}$ can be seen in the bottom left image that results from hardware imperfections. Second, images from a $10\times 10\times 5$ voxel grid, for which QWFC is performed on a simulator, see fig. 11 (c).
4 CONCLUSIONS
In this work, QWFC was proposed as a method to realize WFC on a quantum computer for PCG. The idea is to construct a quantum circuit from which valid content instance (that fulfill the prescribed patterns) can be sampled via measurements. In other words, the quantum circuit acts as a special-purpose random generator for content of a desired form that makes use of the intrinsic randomness of quantum physics. This method is not without its challenges. The corresponding quantum circuit becomes larger from both the content and alphabet size and at the same time deeper from the correlations within the content instance distribution, which makes it difficult to evaluate most use cases on current hardware. For this reason, the quantum-classical hybrid method HWFC was proposed that decomposes the problem into smaller sub-problems which can then be executed on the quantum computer with a higher chance of success.
The experimental results have shown the limitations of QWFC and HWFC, but have also proven that at least simple examples can already be implemented on today’s quantum hardware. Furthermore, all experiments were carried out without quantum error correction [17], which can be used to possibly improve the results. The biggest bottleneck of the circuit complexity is the conditional preparation of probability distributions. Investigating possible improvements in this area with regard to specific PCG use cases or alternative preparation approaches (e.g., qGANs [18]) could therefore be a potential direction for future research.
The conceptual advantage of QWFC is that a circuit, once prepared, represents the distribution of content instances. This means that new instances can be generated by simple measurements without any additional algorithmic effort. A shortcoming of QWFC is that only a fixed order of segments is considered instead of a dynamic (e.g., entropy-based) order that depends on the previously selected segments. The realization of such a generalization of the method could also be a point of origin for further studies.
CWFC is a strong method with many areas of application. In this paper, only visual content creation use cases were presented, but QWFC can also be used for non-visual contents like text [9] in analogy to CWFC [19].
In conclusion, the paper aims to provide an initial approach to QWFC that demonstrates its feasibility but leaves room for improvement and raises the question of useful application areas given the limits of NISQ devices.
5 ACKNOWLEDGMENTS

### v2

@@ 2. V2)
where $U_{k}^{\mathrm{ctrl}}(C^{\prime}):=\bigotimes_{(s,v)\in L_{k}(C^{\prime})}\ket{v-1}\bra{v-1}_{s}\in\bigotimes_{l=1}^{k-1}\mathcal{H}_{\sigma_{l}}$ denotes the control operator, and $U_{k}^{\mathrm{init}}(C^{\prime}):=\ket{C^{\prime}}\bra{0}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ the initialization operator with |$\ket{C^{\prime}}_{\sigma_{k}}:=\sum_{\nu\in A}\sqrt{p_{k}^{\mathrm{v,P}}(\nu|\sigma_{k},C^{\prime})}\ket{\nu-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$. The summation in $U_{k}^{\mathrm{ctrl}}(C^{\prime})$ runs over all adjacent segments $L_{k}(C^{\prime}):=\{(s,v)\,|\,(s,v)\in C^{\prime}\,\land\,\exists\,P\in R_{P}\exists\,d\in P_{d}:\alpha^{d}_{\sigma_{k}s}=1\}\subseteq C^{\prime}$, where $R_{P}:=\{P_{1},\dots,P_{m}\}$ stands for the set of patterns in $R\in\mathcal{R}^{\mathrm{P}}$ with $P_{i}\in\mathcal{P}$ for $i\in[1,m]$ and $P_{d}:=\{d_{1},\dots,d_{n}\}$ stands for the set of directions in the pattern $P\in\mathcal{P}$, Equation 3, with $d_{j}\in[1,D]$ for $j\in[1,n]$. Choosing $C^{\prime}$ instead of $L_{k}(C^{\prime})$ in the summation has no effect on $p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$ and thus |$\ket{C^{\prime}}_{\sigma_{k}}$.
In iteration $k$, only the qubits in $\mathcal{Q}_{k-1}^{+}$ with $\mathcal{Q}_{k}^{+}:=\{q\,|\,q=q_{k}^{i}\in\mathcal{Q}_{\sigma_{k}}\,\forall\,i\in[1,q]\,\land\,\sigma_{k}\in L_{k}^{+}\}$ are affected by $U^{\mathrm{ctrl}}(C^{\prime})$ for some $C^{\prime}\in\mathcal{C}_{\sigma}(k)$, where $L_{k}^{+}:=\{s\,|\,\exists\,C^{\prime}\in\mathcal{C}_{\sigma}(k)\,:\,s\in S^{+}(L_{k}(C^{\prime}))\}$ denotes the affected segments for $k=[1,N]$. Here, $\mathcal{Q}_{k}^{+}\subseteq\{\mathcal{Q}_{1},\dots,\mathcal{Q}_{k}\}$ and $L_{k}^{+}\subseteq\{\sigma_{1},\dots,\sigma_{k}\}$, respectively.
The second operator in Equation 11 is the neutral operator
@@
Initially, all qubits $\mathcal{Q}_{k}$ are in the ground state. Therefore, $U_{k}^{\mathrm{init},0}:=\sum_{v\in[2,W]}\ket{v-1}\bra{v-1}_{\sigma_{k}}\in\mathcal{H}_{\sigma_{k}}$ vanishes. Furthermore, $U_{k}^{\mathrm{ctrl},0}(C^{\prime}):=\left[\bigotimes_{l=1}^{k-1}\mathbbm{1}_{\sigma_{l}}-U_{k}^{\mathrm{ctrl}}(C^{\prime})\right]\otimes\mathbbm{1}_{\sigma_{k}}\in\bigotimes_{l=1}^{k}\mathcal{H}_{\sigma_{l}}$ does not change the state by construction.
@@
To realize $U_{k}$ with quantum gates for a given $k\in[1,N]$, only $U_{k}^{\mathrm{coin}}$ has to be considered (because all other parts of $U_{k}$ have no effect on the resulting state), which consists of a sequence of joint pairs $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, one for each $C^{\prime}\in\mathcal{C}_{\sigma}(k)$. Each of these joint pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{\sigma_{k}}$, which is a well-known task that may, however, require an exponential number of gates [23].
@@
In the original WFC, a pattern conflict leads to a restart of the algorithm. This can also be realized in QWFC by adding an additional “conflict detection qubit” that stores this information, making S1, S2, V1, and V2 obsolete. This optional extension is not further discussed.
Technical Details of HWFC
Presume $H$ partitions such that partition $h\in[1,H]$ contains all of the segment identifiers from the set $S^{h}\subset[1,N]$ with $\bigcup_{h=1}^{H}S^{h}=[1,N]$ and $\bigcap_{h=1}^{H}S^{h}=\{\}$. The already sampled sub-contents $C^{1},\dots,C^{h-1}$ are used as an additional constraint, which has to be taken account in the quantum circuit by replacing $\mathcal{C}_{\sigma}(k)$ in Equations 12 and 13 with $\mathcal{C}^{h}_{\sigma}(k):=\{C\,|\,C=\bigcup_{i=1}^{h-1}C^{i}\cup C^{\prime}\,\forall\,C^{\prime}\in\mathcal{C}_{\sigma}(k)\}$.
ACKNOWLEDGMENTS